OpenStudy (anonymous):

Water freezes at thirty-two degrees Fahrenheit, or zero degrees Celsius. Use temperature conversion formulas to show that these two measures are indeed the same temperature.

OpenStudy (wolf1728):

Here's the formulas:

OpenStudy (wolf1728):

Using the first formula: Fahrenheit = [9/5 * 0] +32 Fahrenheit = 0 + 32 Fahrenheit = 32 Second Formula Celsius = [32-32]*5/9 Celsius = 0 * 5/9 Celsius = 0

OpenStudy (anonymous):

So they're the same?

OpenStudy (wolf1728):

Yes they are the same 32 degrees Fahrenheit = 0 degrees Celsius

OpenStudy (anonymous):

So, what would I write down?

OpenStudy (wolf1728):

Well you could write down the two temperature conversions I did and that would show both those temperatures equal each other.

OpenStudy (anonymous):

Okay, thank you very much!

OpenStudy (wolf1728):

u r welcome